# Scaling Population-Based Reinforcement Learning with GPU Accelerated Simulation

### v3

@@ III-B3 PBRL-DDPG
In DDPG, exploration noise is added to the output of the deterministic actor. As mentioned in Section II-A, different noise levels are added for different environments uniformly within the range . Both these parameters are crucial in controlling the amount of exploration in DDPG agents. In PBRL-DDPG, the hyperparameters optimized during training include the minimum and the maximum bounds for noise levels, i.e., , and the LRs of the actor and the critic. As in PBRL-SAC, the maximum population size in PBRL-DDPG is set to 8 due to the presence of independent replay buffers and GPU memory limitations. Figure 5 shows that PBRL-DDPG achieves significantly better training performance than DDPG on all 4 benchmark tasks.
III-B4 Mutation Comparison
@@
Figure 6 shows the results using 3 different mutation schemes for PBRL-PPO and PBRL-DDPG. As mentioned in Section II-B, the hyperparameters for under-performing agents are generated either by sampling from an original prior distribution, by perturbing the parent’s values through perturbation factors given in Table III, or through the DexPBT mutation scheme presented in [27]. In the latter, the hyperparameters have a probability of getting multiplied or divided by a random number sampled from the uniform distribution, . The results show that the perturbed agents either exceed or are on par with the performance of other mutation schemes in 6 out of 8 experiments. The DexPBT mutation scheme performs better with PBRL-DDPG on Humanoid and Franka Nut Pick tasks, which are less challenging compared to others. The combination of two mutation schemes might discover better exploration strategies for a wider range of tasks. Analyzing the potential synergies between the two remains a prospect for future investigation.
@@
III-C Sim-to-Real Transfer
In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation, executing the actions with PLAI [10]. The robot detects the nuts utilizing Mask-RCNN [32], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10]. Compared to the original task [4], we applied the following changes to make the simulated environment resemble real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to reflect the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose. The parameters used in the simulated environment and the real controller are reported in Table IV.
During experiments, the following policies were deployed, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO. With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase. As shown in Figure 7, PBRL-PPO with has the highest success rate. Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs. ).
Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Figure 8), yet with different success rates, as summarized in Table V. Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation. Unlike the baseline PPO agent, which continued to produce small movements after reaching the target, PBRL-PPO agents remained more stable, leading to a higher success rate. This demonstrates that PBRL agents, while achieving similar rewards to a single agent, learn behaviors that exhibit greater robustness to environment variability due to the diversity in agent populations. Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
@@ III-C Sim-to-Real Transfer
III-D Discussion
While the PBRL agents perform better than the non-evolutionary agents in almost all the experiments, the impact of population size across RL algorithms and tasks shows no consistent pattern. One may hypothesize that larger and more diverse populations might lead to a better final performance. However, the results in this work indicate that using a larger population size does not necessarily yield substantial benefits for every task. This is in contrast to the common belief that population-based methods rely on larger population sizes to effectively explore the hyperparameter space [26, 35]. The optimal population size, instead, depends on various factors, including task complexity, RL algorithm, and interaction dynamics among agents. While larger populations offer increased exploration potential, they also suffer from diminished exploitation capabilities due to increased competition, leading to lower performance in less challenging tasks where smaller populations suffice. Larger population sizes seem to perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
Additionally, the performance of PBRL may be lower than non-evolutionary agents on relatively simpler tasks where optimal hyperparameters are known a priori. This can be noticed on a Humanoid task trained with SAC in Figure 4: indeed, baseline policies achieve a higher reward than PBRL-SAC with 4 agents; nevertheless, 8 agents are capable of outperforming the baseline. Thus, the benefits provided by PBRL will become more apparent for new tasks where ideal hyperparameter ranges are not known in advance. In this sense, PBRL can be thought of as an exploratory approach to search for unknown optimal configurations of newly designed tasks.
@@ IV Conclusion
IV-A Summary
In this paper, a PBRL framework has been employed to train a population of RL agents by making use of high-throughput GPU-accelerated simulation. The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of locomotion and manipulation benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes. The results showed the effectiveness of PBRL in improving final performance through online adaptation of hyperparameters. PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning. Finally, we released the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms.
IV-B Future Works
Many interesting directions remain for future research. An immediate extension could be to use a dedicated fitness metric for each population sub-group to prioritize long-term performance [36]. This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training. Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation or assembly) in the real world with sim-to-real techniques. With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable.

### v4

@@ III-A3 PBRL-DDPG
In DDPG, exploration noise is added to the output of the deterministic actor. As mentioned in Section II-A, different noise levels are added for different environments uniformly within the range . Both these parameters are crucial in controlling the amount of exploration in DDPG agents. In PBRL-DDPG, the hyperparameters optimized during training include the minimum and the maximum bounds for noise levels, i.e., , and the LRs of the actor and the critic. As in PBRL-SAC, the maximum population size in PBRL-DDPG is set to 8 due to the presence of independent replay buffers and GPU memory limitations. Fig. 3 shows that PBRL-DDPG achieves significantly better training performance than DDPG on all 4 benchmark tasks.
III-A4 Mutation Comparison
@@
Fig. 4 shows the results using 3 different mutation schemes for PBRL-PPO and PBRL-DDPG. As mentioned in Section II-B, the hyperparameters for under-performing agents are generated either by randomly sampling from their range (see Tables I–II), by perturbing the parent’s values through perturbation factors given in Table III, or through the DexPBT mutation scheme [27]. In the latter, the hyperparameters have a probability of getting multiplied or divided by a random number sampled from the uniform distribution, . The results manifest that the perturbed agents either exceed or are on par with the performance of other mutation schemes in 6 out of 8 experiments, with DexPBT mutation scheme performs better with PBRL-DDPG on Humanoid and Franka Nut Pick tasks, which are less challenging compared to others. Generally speaking, the perturbation scheme yields less variance (i.e., a more stable learning) compared to DexPBT.
III-B Discussion
Although PBRL agents outperform non-evolutionary agents in nearly all experiments, the influence of population size across different RL algorithms and tasks does not follow a clear or consistent pattern. One might intuitively hypothesize that larger, more diverse populations would improve final performance by enabling broader exploration. However, our results suggest that increasing the population size does not universally enhance agents training. This challenges the common assumption that population-based methods inherently benefit from large populations due to their ability of performing a more thorough exploration of the hyperparameter space [26, 37]. Instead, the optimal population size appears to be task- and algorithm-dependent, influenced by factors such as task complexity and the nature of agents interaction. For simpler tasks, smaller populations may suffice to reach high performance, and additional agents may offer diminishing returns or increase noise during selection. While our evolutionary mechanism keeps the degree of exploitation by propagating successful hyperparameters, larger populations may dilute this effect by introducing more diverse, potentially suboptimal, behavior patterns. In contrast, tasks with gradually increasing complexity — such as Anymal Terrain, which uses curriculum learning — may benefit more from the exploration enabled by larger populations.
Additionally, PBRL performance may lag behind non-evolutionary agents on relatively simpler tasks where optimal hyperparameters are known a priori. This can be noticed on a Humanoid task trained with SAC in Fig. 3: indeed, baseline policies achieve a higher reward than PBRL-SAC with 4 agents; nevertheless, 8 agents are capable of outperforming the baseline. Thus, the benefits provided by PBRL will become more apparent for novel tasks lacking known ideal hyperparameter ranges. In this sense, PBRL can be thought of as an exploratory approach to search for unknown optimal configurations of newly designed tasks.
III-C Sim-to-Real Transfer
For real-world experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation on a physical robot, deploying the actions with PLAI [10]. The robot detects the nuts utilizing Mask-RCNN [38], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10].
@@ III-C Sim-to-Real Transfer
Compared to the original task [4], we applied the following changes to make the simulated environment as close as possible to the real setup: (i) employing a Task-Space Impedance (TSI) controller [39] instead of an Operational-Space motion Controller (OSC) [40] to comply with the actual low-level controller444The control laws are specified in [4] and in reference works [39, 40]; (ii) randomizing the nut’s initial position to resemble the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose. The parameters used in the simulated environment and the real controller are reported in Table IV.
@@
During experiments, we deployed the following policies, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO. With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase. As shown in Fig. 5, PBRL-PPO with has the highest success rate. Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs. ), also highlighting lower variance compared to both baseline PPO and PBRL-PPO with .
@@
Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Fig. 6), yet with different success rates, as summarized in Table V. Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation. This suggests that a better exploration across the agents, favored by PBRL, not only leads to higher rewards, but is also one of the aspects that could lead to a successful real-world transfer [31]. Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
@@ IV Conclusion
In this paper, a PBRL framework has been employed for training a population of RL agents by making use of high-throughput GPU-accelerated simulation. The first simulation results of PBRL using on- and off-policy methods are provided on a series of benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes. The results showed the effectiveness of PBRL in improving final performance by adapting the hyperparameters online. PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning. Finally, we released the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms on challenging robotic tasks.
Many interesting directions remain for future research. An immediate extension could be to use a dedicated fitness metric for each population sub-group to prioritize long-term performance [41]. This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training.
Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation, and assembly) in the real world with sim-to-real techniques. With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable, accounting for diverse skills labeling and learning.